OpenStudy (kellygirl34):

Tablets are on sale for 15% off the original price (t), which can be expressed with the function p(t) = 0.85t. Local taxes are an additional 8% of the discounted price (p), which can be expressed with the function c(p) = 1.08p. Using this information, which of the following represents the final price of a tablet with the discount and taxes applied based on its original price? (2 points) c[p(t)] = 0.918t c(p) + p(t) = 1.93t c(p) ⋅ p(t) = 0.918pt t[c(p)] = 1.93p

OpenStudy (ask_if_im_a_ghost):

c(p)=c(p(t))=c(0.85t)=1.08(0.85t)=0.918t

OpenStudy (kellygirl34):

so what is the answer @Ask_If_Im_a_Ghost

OpenStudy (ask_if_im_a_ghost):

The first one

OpenStudy (kellygirl34):

thats what i thought